Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year: Molding Fabrication Total Machine-hours 25,000 35,000 60,000 Fixed manufacturing overhead costs $ 730,000 $ 280,000 $ 1,010,000 Variable manufacturing overhead cost per machine-hour $ 5.80 $ 5.80 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs: Job D-70: Molding Fabrication Total Direct materials cost $ 373,000 $ 328,000 $ 701,000 Direct labor cost $ 220,000 $ 150,000 $ 370,000 Machine-hours 17,000 8,000 25,000 Job C-200: Molding Fabrication Total Direct materials cost $ 300,000 $ 250,000 $ 550,000 Direct labor cost $ 180,000 $ 280,000 $ 460,000 Machine-hours 8,000 27,000 35,000 Delph had no underapplied or overapplied manufacturing overhead during the year. 2. Assume Delph uses departmental predetermined overhead rates based on

machine-hours. a. Compute the departmental predetermined overhead rates. b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200. c. If Delph establishes bid prices that are 140% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200? d. What is Delph’s cost of goods sold for the year?
Answer #1
2a
Predetermined
overhead rate
Molding Department 35.00 =5.80+(730000/25000)
Fabrication Department 13.80 =5.80+(280000/35000)
b
Job D-70 Job C-200
Direct materials cost 701000 550000
Direct labor cost 370000 460000
Overhead:
Molding Department 595000 280000
Fabrication Department 110400 372600
Total manufacturing cost 1776400 1662600
c
Job D-70 Job C-200
Total manufacturing cost 1776400 1662600
Add: Markup @40% 710560 665040
Bid prices 2486960 2327640
d
Total manufacturing cost Job D-70 1776400
Total manufacturing cost Job C-200 1662600
Cost of goods sold 3439000
Workings:
Overhead applied:
Job D-70 Job C-200
Molding Department =17000*35 =8000*35
Fabrication Department =8000*13.80 =27000*13.80